OpenStudy (anonymous):

CAN SOME 1 EXPLAIN THE EQUATION FOR SLOPE

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

pooja195 (pooja195):

Point slope?

OpenStudy (anonymous):

APPERICIATE

pooja195 (pooja195):

\[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

pooja195 (pooja195):

Or slope slope?

OpenStudy (anonymous):

ITS FINE FIGURED IT OUT

pooja195 (pooja195):

._. ok welll welcome to openstudy spongebob :P

OpenStudy (anonymous):

THANKS I FAN U

pooja195 (pooja195):

:3

OpenStudy (anonymous):

BE MY FAN